A. In your hypothetical economy, The Gross Domestic Product at Current Prices for 2021 was estimated at K10000 compared to K8000 recorded in 2020. In addition, Gross Domestic Product at constant market prices was K8000 in 2021 and K600 in 2020. Calculate the GDP deflator for 2020 and 2021 and the inflation rate between the two years

B. Explain how strong trade unions can cause unemployment and inflation in an economy.


1
Expert's answer
2021-09-06T07:05:13-0400

A. 

  • Nominal GDP (2021) = K10000
  • Nominal GDP (2020) = K8000
  • Real GDP (2021) = K8000
  • Real GDP (2020) = K600

The GDP deflator for 2020

"GDP\, deflator = \frac{Nominal\, GDP}{Real\, GDP}\, \times 100\\
GDP\, deflator = \frac{8000}{600}\, \times 100\\
GDP\, deflator = 1333.33"

The GDP deflator for 2021

"GDP\, deflator = \frac{Nominal\, GDP}{Real\, GDP}\, \times 100\\
GDP\, deflator = \frac{10000}{8000}\, \times 100\\
GDP\, deflator = 125"

The inflation rate between the two years

"\mathbf{Inflation\,rate\, = \frac{GDP\, Deflator\, in\, 2021\, -\,GDP\, Deflator\, in\, 2020 }{GDP\, Deflator\, in\, 2020}\times 100 }\\
\mathbf{Inflation\,rate\, = \frac{125\, -\,1333.33 }{1333.33}\times 100}\\
\mathbf{Inflation\,rate\, = -90.62\%}"

Note: Negative inflation shows that between 2020 and 2021, there was deflation.

B.

Strong trade unions are likely to succeed in bargaining for higher wages for employees. Higher wages that exceed the equilibrium wages will increase the cost of production for producers. Therefore, the producers will reduce the number of employees to reduce the costs of production. This situation will cause a rise in unemployment. Furthermore, the increased cost of production resulting from higher wages will cause a reduction in the short-run aggregate supply. The fall in the short-run aggregate supply curve will cause an increase in inflation and a decline in real output.